OpenStudy (anonymous):

equivalent to what is a expression equivalent to log18 – log(p + 2)?

Nnesha (nnesha):

familiar with the log properties ?

OpenStudy (anonymous):

nall but they gave me choices log p+2/18 , log 18/p+2 , log20/p log{18 times (p+2)]

Nnesha (nnesha):

nope log property that suppose to be on ur notes :-)

Nnesha (nnesha):

anyways quotient rule\[\large\rm log_b y - \log_b x = \log_b \frac{ x }{ y}\]to condense you can change subtraction to division product rule \[\large\rm log_b x + \log_b y = \log_b( x \times y )\] addition ----> multiplication power rule \[\large\rm log_b x^y = y \log_b x\] hint: \[\large\rm log_b x = \log_b y\] \[\large\rm\cancel { log_b} x = \cancel{\log_b} y\] x=y

Nnesha (nnesha):

so which one you should apply ?

OpenStudy (anonymous):

the second one

OpenStudy (anonymous):

so my answer is D

Nnesha (nnesha):

\[\huge\rm log18 \color{ReD}{–} \log(p + 2)\] there is a negative sign so which property you should apply here ?

OpenStudy (anonymous):

the quotient rule

Nnesha (nnesha):

yes that's right so \[\huge\rm log18 \color{ReD}{–} \log(p + 2)= ???\]

OpenStudy (anonymous):

log{18 times (p+2)]

Nnesha (nnesha):

quotient \[\large\rm log_b x- \log_b y = \log_b \frac{ x }{ y}\] same like this property log x/y

Nnesha (nnesha):

quotient rule when there is a negative sign change that to DIVISION

OpenStudy (anonymous):

oh so its log 18/p+2 ,

Nnesha (nnesha):

yes right

OpenStudy (anonymous):

i got lot more questions help me